# The Complexity of Sequential Prediction in Dynamical Systems

### v1

@@ Appendix A Combinatorial dimensions
Note $|\mathcal{T}^{x}_{t}(\varepsilon_{\leq t})|=|\mathcal{T}^{x}_{0}|+t$ for all $t\in[T]$. Moreover, for every path $\varepsilon\in\{-1,1\}^{T}$, there exists a function $\sigma\in\{-1,1\}^{\mathbbm{N}\cup\{0\}}$ such that $\sigma_{|\mathcal{T}^{x}_{0}|}=\varepsilon_{1}$ and $\sigma_{|\mathcal{T}^{x}_{t}(\varepsilon_{\leq t})|}=\varepsilon_{t+1}$ for all $t\in[T-1].$ Thus, the function $f_{\sigma}\in\mathcal{F}_{S}$ shatters the path $\varepsilon$ and the tree $\mathcal{T}^{x}$ is shattered by $\mathcal{F}_{S}$. We claim that $\texttt{B}(\mathcal{T}^{x})=|S\cap\{|\mathcal{T}^{x}_{0}|,...,|\mathcal{T}^{x}_{0}|+T-1\}.$ To see this, observe that for every path $\varepsilon\in\{-1,1\}^{T}$, we have
@@
Thus, $\texttt{B}(\mathcal{T}^{x})=|S\cap\{|x_{0}|,...,|x_{0}|+T-1\}|$ and $\texttt{C}_{T}(\mathcal{F})\geq\sup_{x\in\mathbb{Z}}\texttt{B}(\mathcal{T}^{x})=\sup_{x\in\mathbb{Z}}|S\cap\{|x|,...,|x|+T-1\}|.$ This completes the proof.
Appendix E Proof of Theorem 3.4
The proof of (i) follows from Theorem 3.1 and the fact that $\texttt{C}_{T}(\mathcal{F})\leq\texttt{dim}(\mathcal{F})$. To prove (ii), observe that by Theorem 3.1, it suffices to show that when $\texttt{dim}(\mathcal{F})=\infty$, we have that $\texttt{C}_{T}(\mathcal{F})=\omega(1)$. If $\texttt{dim}(\mathcal{F})=\infty$, then for every $d\in\mathbb{N}$, there exists a shattered trajectory tree $\mathcal{T}$ such that $\texttt{B}(\mathcal{T})>d$. Thus, for every $d\in\mathbb{N}$, there exists a depth $d^{\prime}\in\mathbb{N}$, such that for every $T\geq d^{\prime}$, there exists a shattered tree $\mathcal{T}$ of depth $T$ with $\texttt{B}(\mathcal{T})>d$. In other words, for every $d\in\mathbb{N}$, there exists a $d^{\prime}\in\mathbb{N}$ such that for all $T\geq d^{\prime}$ we have that $\texttt{C}_{T}(\mathcal{F})>d$. By definition of $\omega(\cdot)$, this means that $\texttt{C}_{T}(\mathcal{F})=\omega(1)$ as needed.
Appendix F Proof of Theorem 3.5
Proof.
(of (i) in Theorem 3.5) To prove (i), pick $S=\{2^{t}:t\in\mathbb{N}\cup\{0\}\}$ and consider the function class $\mathcal{F}_{S}$ from Theorem 3.3. It is not too hard to see that $\texttt{C}_{T}(\mathcal{F})=\Theta(\log(T))$. In addition, one can verify that every finite subset of $S$ is a shattered set according to Definition 9. Indeed, consider any finite subset $A\subset S$. For every $i\in[|A|]$, let $A_{i}$ denote the $i$’th element of $A$ after sorting $A$ in increasing order. Then, observe that by the construction of $\mathcal{F}_{S}$, for every sequence $\varepsilon\in\{-1,1\}^{|A|}$, there exists a function $f_{\varepsilon}\in\mathcal{F}_{S}$ such that $f_{\varepsilon}(A_{i})=\varepsilon_{i}(A_{i}+1)$ for every $i\in[|A|].$ By letting =$\mathcal{H}=\{f_{\varepsilon}:\varepsilon\in\{-1,1\}^{|A|}\}$ in Definition 9, one can verify that $|\mathcal{H}|=2^{|A|}<\infty$, and for every $x\in A$ and $h\in\mathcal{H}$, there exists a $g\in\mathcal{H}$ such that $h(x)=-g(x)$ and $h(z)=g(z)$ for all $z\in A\setminus\{x\}$. Thus, $\mathcal{F}_{S}$ shatters $A\subset S$. Since finite subsets of $S$ can be arbitrary large, we have that $\texttt{DS}(\mathcal{F})=\infty$. ∎
Proof.
@@
Consider the function class =$\mathcal{F}=\{f_{\sigma}:\sigma\in\{-1,1\}^{\mathbb{N}}\}.$ We now claim that $\texttt{E}_{1}(\mathcal{F})=\infty$. To see this, fix a depth $d\in\mathbb{N}$, and consider the following trajectory tree $\mathcal{T}$ of depth $d$. Let the root node be labeled by $1$, that is $\mathcal{T}_{0}=1$. For all $t\in[d]$ and $\varepsilon\in\{-1,1\}^{t}$, let $\mathcal{T}_{t}(\varepsilon)=a_{\tilde{\varepsilon}}(t+1)$ where $\tilde{\varepsilon}$ denotes an arbitrary extension of $\varepsilon$ over $\mathbb{N}$. Note that the completion can be arbitrary because the value of $a_{\sigma}(t+1)$ for any $\sigma\in\{-1,1\}^{\mathbb{N}}$ depends only on the prefix $(\sigma_{1},...,\sigma_{t})$. One can verify that such a tree $\mathcal{T}$ is a complete binary tree of depth $d$ where the root node is labeled with $1$ and the internal nodes on depth $i\geq 1$ are labeled from left to right by $12^{i},2^{i}+1,...,2^{i+1}-$ 1. Thus, it is clear that $\texttt{B}(\mathcal{T})=d$ since for every internal node including the root, its two children are labeled by differing states. In addition, observe that for every path $\varepsilon\in\{-1,1\}^{d}$ down $\mathcal{T}$, the function $f_{\tilde{\varepsilon}}\in\mathcal{F}$ shatters the associated sequence of states, where $\tilde{\varepsilon}$ again is an arbitrary completion of $\varepsilon$ over $\mathbb{N}$. Indeed, fix a $\varepsilon\in\{-1,1\}^{d}$, a completion $\tilde{\varepsilon}\in\{-1,1\}^{\mathbb{N}}$, and consider any $t\in[d]$. Then, by definition of $\mathcal{T}$, we have that $\mathcal{T}_{t-1}(\varepsilon_{<t})=a_{\tilde{\varepsilon}}(t)$ and $\mathcal{T}_{t}(\varepsilon_{\leq t})=a_{\tilde{\varepsilon}}(t+1)$. Consider the function $f_{\tilde{\varepsilon}}\in\mathcal{F}$. By definition of $f_{\tilde{\varepsilon}}$, we have that $f_{\tilde{\varepsilon}}(a_{\tilde{\varepsilon}}(t))=a_{\tilde{\varepsilon}}(t+1)$, which implies that $f_{\tilde{\varepsilon}}(\mathcal{T}_{t-1}(\varepsilon_{<t}))=\mathcal{T}_{t}(\varepsilon_{\leq t})$. Since $t\in[d]$ was arbitrary, this is true for every $t\in[d]$, and thus $f_{\tilde{\varepsilon}}$ shatters the path $\varepsilon$ down $\mathcal{T}$ as claimed. Since $\varepsilon\in\{-1,1\}^{d}$ was also arbitrary, we have that the entire tree $\mathcal{T}$ is shattered by $\mathcal{F}$. Finally, since $d\in\mathbb{N}$ was arbitrary, this is true for arbitrarily large depths. Thus, $\texttt{E}_{1}(\mathcal{F})=\infty$.
@@
We now show that $\texttt{DS}(\mathcal{F})=1$ by proving that $\mathcal{F}$ cannot DS-shatter any two instances $x_{1},x_{2}\in\mathcal{X}$. Our proof will be in cases. First, observe that if either $x_{1}=\star$ or $x_{2}=\star$, then $(x_{1},x_{2})$ cannot be shattered since all functions in $\mathcal{F}$ will output $\star$ on either $x_{1}$ or $x_{2}$. Thus, without loss of generality, suppose both $x_{1},x_{2}\in\mathbb{N}$ and $x_{1}<x_{2}$. Consider any finite subset $\mathcal{H}\subset\mathcal{F}$ and suppose there exists a function $h\in\mathcal{H}$ such that $h(x_{2})\neq\star$. Then, in order to shatter $(x_{1},x_{2})$, there must exist a function $g\in\mathcal{H}$ such that $h(x_{1})\neq g(x_{1})$ but $h(x_{2})=g(x_{2})$. However, if $h(x_{2})=g(x_{2})\neq\star$, then it must be the case that $h(x_{1})=g(x_{1})$. To see why, fix an instance $x\in\mathbb{N}$, and suppose $f_{\sigma_{1}}(x)=f_{\sigma_{2}}(x)\neq\star$. Then, by properties (2) and (4) above, it must be the case that $a^{-1}_{\sigma_{1}}(x)=a^{-1}_{\sigma_{2}}(x)=c$ and $a_{\sigma_{1}}(i)=a_{\sigma_{2}}(i)$ for all $i\leq c$. Thus, if $x_{1}<x_{2}$ and $f_{\sigma_{1}}(x_{2})=f_{\sigma_{2}}(x_{2})\neq\star$, we must have that $f_{\sigma_{1}}(x_{1})=f_{\sigma_{2}}(x_{1})$ because either $x_{1}\notin\text{im}(a_{\sigma_{1}})\cup\text{im}(a_{\sigma_{2}})$ or $a^{-1}_{\sigma_{1}}(x_{1})<a^{-1}_{\sigma_{1}}(x_{2})=a^{-1}_{\sigma_{2}}(x_{2})$. Thus, if $\mathcal{H}$ were to satisfy the property in Definition 9, there cannot exist a hypothesis $h\in\mathcal{H}$ such that $h(x_{2})\neq\star$. However, if for every $h\in\mathcal{H}$, we have that $h(x_{2})=\star$, then for every $h\in\mathcal{H}$, there cannot exist a $g\in\mathcal{H}$ such that $h(x_{1})=g(x_{1})$ and $h(x_{2})\neq g(x_{2})$. Therefore, the two points $(x_{1},x_{2})$ cannot be shattered. Since $(x_{1},x_{2})$ and $\mathcal{H}\subset\mathcal{F}$ were arbitrary, this is true for all such points, implying that $\texttt{DS}(\mathcal{F})\leq 1$. Since $|\mathcal{F}|\geq 2$, we have also that $\texttt{DS}(\mathcal{F})\geq 1$, completing the proof that $\texttt{DS}(\mathcal{F})=1$. ∎
Proof.
@@
(of (iii) in Theorem 3.5) To prove part (iii) of Theorem 3.5, we reduce learning dynamical systems to online multiclass classification. Let $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ and $\mathcal{B}$ be any (potentially randomized) online learner for $\mathcal{F}$ for online multiclass classification with expected regret bound $R$. We will construct a learner $\mathcal{A}$ that uses $\mathcal{B}$ as a subroutine such that $\texttt{M}_{\mathcal{A}}(T,\mathcal{F})\leq R$.
@@
for $t=1,...,T$ do
       Pass $x_{t-1}$ to $\mathcal{B}$ and receive prediction $\hat{z}_{t}.$ Predict $\hat{x}_{t}=\hat{z}_{t}$. Receive next state $x_{t}$ and update $\mathcal{B}$ using labeled instance $(x_{t-1},x_{t}).$
end for
@@
where (i) follows from the expected regret guarantee of $\mathcal{B}$ and (ii) follows from the fact that the stream of states is realizable. Thus, $\inf_{\mathcal{A}}\texttt{M}_{\mathcal{A}}(T,\mathcal{F})\leq R$. Since $\mathcal{B}$ is always guaranteed to observe a realizable sequence of labeled instances, picking the SOA [Littlestone, 1987, Daniely et al., 2011] gives that $R=\texttt{L}(\mathcal{F})$. Since the SOA is deterministic, this choice of $\mathcal{B}$ implies that $\mathcal{A}$ is deterministic. Part (i) then follows from that the fact that for any deterministic learner $\mathcal{A}$ we have $\texttt{M}_{\mathcal{A}}(T,\mathcal{F})\geq\texttt{C}_{T}(\mathcal{F})$. ∎
Proof.
@@
(of (iv) in Theorem 3.5) To prove part (iv), let $\mathcal{X}=[0,1]$ and consider the class of thresholds =$\mathcal{F}=\{x\mapsto\mathbbm{1}\{x\geq a\}:a\in(0,1)\}$. It is well known that $\texttt{L}(\mathcal{F})=\infty$. On the other hand, note that $\dim(\mathcal{F})=1$ since $\mathcal{F}(1)=\{1\}$, $\mathcal{F}(0)=\{0\}$, and $\mathcal{F}(x)=\{0,1\}$ for all $x\in(0,1)$. ∎
Appendix G Discrete Linear Systems and Linear Boolean Networks
G.1 Discrete Linear Systems
Proof.

### v2

@@ Appendix A Proofs for Realizable learnability
Note $|\mathcal{T}^{x}_{t}(\varepsilon_{\leq t})|=|\mathcal{T}^{x}_{0}|+t$ for all $t\in[T]$. Moreover, for every path $\varepsilon\in\{-1,1\}^{T}$, there exists a function $\sigma\in\{-1,1\}^{\mathbbm{N}\cup\{0\}}$ such that $\sigma_{|\mathcal{T}^{x}_{0}|}=\varepsilon_{1}$ and $\sigma_{|\mathcal{T}^{x}_{t}(\varepsilon_{\leq t})|}=\varepsilon_{t+1}$ for all $t\in[T-1].$ Thus, the function $f_{\sigma}\in\mathcal{F}_{S}$ shatters the path $\varepsilon$ and the tree $\mathcal{T}^{x}$ is shattered by $\mathcal{F}_{S}$. We claim that $\operatorname{B}(\mathcal{T}^{x})=|S\cap\{|\mathcal{T}^{x}_{0}|,...,|\mathcal{T}^{x}_{0}|+T-1\}.$ To see this, observe that for every path $\varepsilon\in\{-1,1\}^{T}$, we have
@@
Thus, $\operatorname{B}(\mathcal{T}^{x})=|S\cap\{|x_{0}|,...,|x_{0}|+T-1\}|$ and $\operatorname{C}_{T}(\mathcal{F})\geq\sup_{x\in\mathbb{Z}}\operatorname{B}(\mathcal{T}^{x})=\sup_{x\in\mathbb{Z}}|S\cap\{|x|,...,|x|+T-1\}|.$ This completes the proof.
Appendix C Proof of Theorem 3.4
The proof of (i) follows from Theorem 3.1 and the fact that $\operatorname{C}_{T}(\mathcal{F})\leq\operatorname{Bd}(\mathcal{F})$. To prove (ii), observe that by Theorem 3.1, it suffices to show that when $\operatorname{Bd}(\mathcal{F})=\infty$, we have that $\operatorname{C}_{T}(\mathcal{F})=\omega(1)$. If $\operatorname{Bd}(\mathcal{F})=\infty$, then for every $d\in\mathbb{N}$, there exists a shattered trajectory tree $\mathcal{T}$ such that $\operatorname{B}(\mathcal{T})>d$. Thus, for every $d\in\mathbb{N}$, there exists a depth $d^{\prime}\in\mathbb{N}$, such that for every $T\geq d^{\prime}$, there exists a shattered tree $\mathcal{T}$ of depth $T$ with $\operatorname{B}(\mathcal{T})>d$. In other words, for every $d\in\mathbb{N}$, there exists a $d^{\prime}\in\mathbb{N}$ such that for all $T\geq d^{\prime}$ we have that $\operatorname{C}_{T}(\mathcal{F})>d$. By definition of $\omega(\cdot)$, this means that $\operatorname{C}_{T}(\mathcal{F})=\omega(1)$ as needed.
Appendix D Proof of Theorem 3.5
Proof D.1.
(of (i) in Theorem 3.5) To prove (i), pick $S=\{2^{t}:t\in\mathbb{N}\cup\{0\}\}$ and consider the function class $\mathcal{F}_{S}$ from Theorem 3.2. It is not too hard to see that $\operatorname{C}_{T}(\mathcal{F})=\Theta(\log(T))$. In addition, one can verify that every finite subset of $S$ is a shattered set according to Definition 2.7. Indeed, consider any finite subset $A\subset S$. For every $i\in[|A|]$, let $A_{i}$ denote the $i$’th element of $A$ after sorting $A$ in increasing order. Then, observe that by the construction of $\mathcal{F}_{S}$, for every sequence $\varepsilon\in\{-1,1\}^{|A|}$, there exists a function $f_{\varepsilon}\in\mathcal{F}_{S}$ such that $f_{\varepsilon}(A_{i})=\varepsilon_{i}(A_{i}+1)$ for every $i\in[|A|].$ By letting =$\mathcal{H}=\{f_{\varepsilon}:\varepsilon\in\{-1,1\}^{|A|}\}$ in Definition 2.7, one can verify that $|\mathcal{H}|=2^{|A|}<\infty$, and for every $x\in A$ and $h\in\mathcal{H}$, there exists a $g\in\mathcal{H}$ such that $h(x)=-g(x)$ and $h(z)=g(z)$ for all $z\in A\setminus\{x\}$. Thus, $\mathcal{F}_{S}$ shatters $A\subset S$. Since finite subsets of $S$ can be arbitrary large, we have that $\operatorname{DS}(\mathcal{F})=\infty$.
Proof D.2.
@@
Consider the function class =$\mathcal{F}=\{f_{\sigma}:\sigma\in\{-1,1\}^{\mathbb{N}}\}.$ We now claim that $\operatorname{C}_{T}(\mathcal{F})=T$. To see this, fix a depth $d\in\mathbb{N}$, and consider the following trajectory tree $\mathcal{T}$ of depth $d$. Let the root node be labeled by $1$, that is $\mathcal{T}_{0}=1$. For all $t\in[d]$ and $\varepsilon\in\{-1,1\}^{t}$, let $\mathcal{T}_{t}(\varepsilon)=a_{\tilde{\varepsilon}}(t+1)$ where $\tilde{\varepsilon}$ denotes an arbitrary extension of $\varepsilon$ over $\mathbb{N}$. Note that the completion can be arbitrary because the value of $a_{\sigma}(t+1)$ for any $\sigma\in\{-1,1\}^{\mathbb{N}}$ depends only on the prefix $(\sigma_{1},...,\sigma_{t})$. One can verify that such a tree $\mathcal{T}$ is a complete binary tree of depth $d$ where the root node is labeled with $1$ and the internal nodes on depth $i\geq 1$ are labeled from left to right by $12^{i},2^{i}+1,...,2^{i+1}-$ 1. Thus, it is clear that $\operatorname{B}(\mathcal{T})=d$ since for every internal node including the root, its two children are labeled by differing states. In addition, observe that for every path $\varepsilon\in\{-1,1\}^{d}$ down $\mathcal{T}$, the function $f_{\tilde{\varepsilon}}\in\mathcal{F}$ shatters the associated sequence of states, where $\tilde{\varepsilon}$ again is an arbitrary completion of $\varepsilon$ over $\mathbb{N}$. Indeed, fix a $\varepsilon\in\{-1,1\}^{d}$, a completion $\tilde{\varepsilon}\in\{-1,1\}^{\mathbb{N}}$, and consider any $t\in[d]$. Then, by definition of $\mathcal{T}$, we have that $\mathcal{T}_{t-1}(\varepsilon_{<t})=a_{\tilde{\varepsilon}}(t)$ and $\mathcal{T}_{t}(\varepsilon_{\leq t})=a_{\tilde{\varepsilon}}(t+1)$. Consider the function $f_{\tilde{\varepsilon}}\in\mathcal{F}$. By definition of $f_{\tilde{\varepsilon}}$, we have that $f_{\tilde{\varepsilon}}(a_{\tilde{\varepsilon}}(t))=a_{\tilde{\varepsilon}}(t+1)$, which implies that $f_{\tilde{\varepsilon}}(\mathcal{T}_{t-1}(\varepsilon_{<t}))=\mathcal{T}_{t}(\varepsilon_{\leq t})$. Since $t\in[d]$ was arbitrary, this is true for every $t\in[d]$, and thus $f_{\tilde{\varepsilon}}$ shatters the path $\varepsilon$ down $\mathcal{T}$ as claimed. Since $\varepsilon\in\{-1,1\}^{d}$ was also arbitrary, we have that the entire tree $\mathcal{T}$ is shattered by $\mathcal{F}$. Finally, since $d\in\mathbb{N}$ was arbitrary, this is true for arbitrarily large depths. Thus, $\operatorname{C}_{T}(\mathcal{F})=T$.
@@
We now show that $\operatorname{DS}(\mathcal{F})=1$ by proving that $\mathcal{F}$ cannot DS-shatter any two instances $x_{1},x_{2}\in\mathcal{X}$. Our proof will be in cases. First, observe that if either $x_{1}=\star$ or $x_{2}=\star$, then $(x_{1},x_{2})$ cannot be shattered since all functions in $\mathcal{F}$ will output $\star$ on either $x_{1}$ or $x_{2}$. Thus, without loss of generality, suppose both $x_{1},x_{2}\in\mathbb{N}$ and $x_{1}<x_{2}$. Consider any finite subset $\mathcal{H}\subset\mathcal{F}$ and suppose there exists a function $h\in\mathcal{H}$ such that $h(x_{2})\neq\star$. Then, in order to shatter $(x_{1},x_{2})$, there must exist a function $g\in\mathcal{H}$ such that $h(x_{1})\neq g(x_{1})$ but $h(x_{2})=g(x_{2})$. However, if $h(x_{2})=g(x_{2})\neq\star$, then it must be the case that $h(x_{1})=g(x_{1})$. To see why, fix an instance $x\in\mathbb{N}$, and suppose $f_{\sigma_{1}}(x)=f_{\sigma_{2}}(x)\neq\star$. Then, by properties (2) and (4) above, it must be the case that $a^{-1}_{\sigma_{1}}(x)=a^{-1}_{\sigma_{2}}(x)=c$ and $a_{\sigma_{1}}(i)=a_{\sigma_{2}}(i)$ for all $i\leq c$. Thus, if $x_{1}<x_{2}$ and $f_{\sigma_{1}}(x_{2})=f_{\sigma_{2}}(x_{2})\neq\star$, we must have that $f_{\sigma_{1}}(x_{1})=f_{\sigma_{2}}(x_{1})$ because either $x_{1}\notin\text{im}(a_{\sigma_{1}})\cup\text{im}(a_{\sigma_{2}})$ or $a^{-1}_{\sigma_{1}}(x_{1})<a^{-1}_{\sigma_{1}}(x_{2})=a^{-1}_{\sigma_{2}}(x_{2})$. Thus, if $\mathcal{H}$ were to satisfy the property in Definition 2.7, there cannot exist a hypothesis $h\in\mathcal{H}$ such that $h(x_{2})\neq\star$. However, if for every $h\in\mathcal{H}$, we have that $h(x_{2})=\star$, then for every $h\in\mathcal{H}$, there cannot exist a $g\in\mathcal{H}$ such that $h(x_{1})=g(x_{1})$ and $h(x_{2})\neq g(x_{2})$. Therefore, the two points $(x_{1},x_{2})$ cannot be shattered. Since $(x_{1},x_{2})$ and $\mathcal{H}\subset\mathcal{F}$ were arbitrary, this is true for all such points, implying that $\operatorname{DS}(\mathcal{F})\leq 1$. Since $|\mathcal{F}|\geq 2$, we have also that $\operatorname{DS}(\mathcal{F})\geq 1$, completing the proof that $\operatorname{DS}(\mathcal{F})=1$.
Proof D.3.
@@
(of (iii) in Theorem 3.5) To prove part (iii) of Theorem 3.5, we reduce learning dynamical systems to online multiclass classification. Let $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ and $\mathcal{B}$ be any (potentially randomized) online learner for $\mathcal{F}$ for online multiclass classification with expected regret bound $R$. We will construct a learner $\mathcal{A}$ that uses $\mathcal{B}$ as a subroutine such that $\operatorname{M}_{\mathcal{A}}(T,\mathcal{F})\leq R$.
@@
1
2for $t=1,...,T$ do
3
4      Pass $x_{t-1}$ to $\mathcal{B}$ and receive prediction $\hat{z}_{t}.$
5      Predict $\hat{x}_{t}=\hat{z}_{t}$.
6      Receive next state $x_{t}$ and update $\mathcal{B}$ using labeled instance $(x_{t-1},x_{t}).$
7 end for
@@
where (i) follows from the expected regret guarantee of $\mathcal{B}$ and (ii) follows from the fact that the stream of states is realizable. Thus, $\inf_{\mathcal{A}}\operatorname{M}_{\mathcal{A}}(T,\mathcal{F})\leq R$. Since $\mathcal{B}$ is always guaranteed to observe a realizable sequence of labeled instances, picking the SOA (Littlestone, 1987; Daniely et al., 2011) gives that $R=\operatorname{L}(\mathcal{F})$. Since the SOA is deterministic, this choice of $\mathcal{B}$ implies that $\mathcal{A}$ is deterministic. Part (i) then follows from that the fact that for any deterministic learner $\mathcal{A}$ we have $\operatorname{M}_{\mathcal{A}}(T,\mathcal{F})\geq\operatorname{C}_{T}(\mathcal{F})$.
Proof D.4.
@@
(of (iv) in Theorem 3.5) To prove part (iv), let $\mathcal{X}=[0,1]$ and consider the class of thresholds =$\mathcal{F}=\{x\mapsto\mathbbm{1}\{x\geq a\}:a\in(0,1)\}$. It is well known that $\operatorname{L}(\mathcal{F})=\infty$. On the other hand, note that $\dim(\mathcal{F})=1$ since $\mathcal{F}(1)=\{1\}$, $\mathcal{F}(0)=\{0\}$, and $\mathcal{F}(x)=\{0,1\}$ for all $x\in(0,1)$.
Appendix E Proofs for Linear Systems
Proof E.1.